OpenStudy (pottersheep):

Plug in the numbers for me please? Formula: d=2rh+h^2) ^1/2 ...h = 200 r: 6370

OpenStudy (pottersheep):

Formula: d=(2rh+h^2) ^1/2 ...h = 200 r: 6370

OpenStudy (pottersheep):

The answer says 1609....but thats not what I get! How about you?

OpenStudy (gw2011):

The answer is 1609. Here is how I got it: d=(2rh+h^2)^1/2 which means you need to take the square root of this whole thing. d=square root [(2)(6370)(200)+(200)^2] d=square root (2,548,000+40,000) d=square root (2,588,000) d=1608.7=1609